OpenStudy (anonymous):

Some polynomial F(x,a) What is this polynomial? After simplification [f(x)-f(a)]/(x-a) = F(x,a)/[(2x^2+2)(2a^2+2)] what is the polynomial of F(x,a)? I got a^2-x^2 which is wrong

OpenStudy (anonymous):

Oh yeah the original function was f(x)=1/ 2x^2+2

OpenStudy (anonymous):

\[\large \frac{(\frac{1}{ 2x^2+2}-\frac{1}{ 2a^2+2})}{x-a}\]Is this your ques?

OpenStudy (anonymous):

I think so and then it gets simplified to F(x,a)/[(2x^2+2)(2a^2+2)] is the polynomial the stuff on top?

OpenStudy (anonymous):

WHAT DOES THIS MEAN?